### v2

# Coordinated Disclosure for AI: Beyond Security Vulnerabilities

Abstract
Harm reporting in the field of Artificial Intelligence (AI) currently operates on an ad hoc basis, lacking a structured process for disclosing or addressing algorithmic flaws. In contrast, the Coordinated Vulnerability Disclosure (CVD) ethos and ecosystem play a pivotal role in software security and transparency. Globally, there are ongoing efforts to establish frameworks that promote transparency and collaboration in addressing AI-related issues, though challenges persist. Algorithmic flaws in machine learning (ML) models present distinct challenges compared to traditional software vulnerabilities, warranting a specialized approach. To address this gap, we propose the implementation of a dedicated Coordinated Flaw Disclosure (CFD) framework tailored to the intricacies of machine learning and artificial intelligence issues. This paper delves into the historical landscape of disclosures in ML, encompassing the ad hoc reporting of harms and the emergence of participatory auditing. By juxtaposing these practices with the well-established disclosure norms in cybersecurity, we argue that the broader adoption of CFD has the potential to enhance public trust through transparent processes that carefully balance the interests of both organizations and the community.
1 Introduction
On December 27, 2023, The New York Times filed a lawsuit against OpenAI and Microsoft for copyright infringement, alleging that ChatGPT was trained on the Times’ published works without authorization, providing evidence that ChatGPT generated articles verbatim (Mac 2023). This action contributed to a growing series of lawsuits against AI providers (O’Brien 2023; Authors Guild 2023; Fenwick & West LLP 2023; ACLU 2022). The specific instances referenced in The New York Times’ lawsuit are challenging to replicate, and as non-parties to the case, we lack insight into whether this is attributable to selective presentation by The New York Times or a true systematic memorization issue in ChatGPT’s design. The Times lawsuit has garnered significant public attention, at least partly due to The New York Times’ financial resources and extensive media reach – factors often unavailable to smaller harmed parties. This highlights the influence of such factors in shaping public discourse, which, in turn, has the potential to improve AI systems via outside pressure (Raji and Buolamwini 2022). Unexpected model behaviors, outside of its intended scope and purpose, underscore the need for robust frameworks to identify and address these issues effectively. This stands in contrast to the Common Vulnerabilities and Exposures (CVE) ecosystem, where vulnerabilities in software systems are systematically cataloged and addressed, often with greater clarity and consensus within the cybersecurity community.
Historically, AI accountability initiatives have relied on periodic audits, emphasizing repetitive assessments but lacking a structured reporting framework for user-identified issues post-deployment, with recent policy efforts acknowledging this issue (European Commission 2020). This audit-centric paradigm is reflected in influential policies such as the U.S. Executive Order on AI (The White House 2023a), the EU’s draft AI Act (European Commission 2021), and New York City’s Local Law 144 (New York City Department of Consumer and Worker Protection 2021). However, this approach falls short when compared to the more comprehensive Coordinated Vulnerability Disclosure (CVD) processes standard in software security.
Coordinated Vulnerability Disclosure (CVD) plays a crucial role as a mechanism for independent researchers to report newly identified vulnerabilities to affected vendors and the public (MultiMedia LLC 1999). This process enables transparent remediation before potential exploitation by malicious actors and has become a vital practice enshrined in government regulations and industry standards. Notably, the FDA mandates the implementation of CVD programs for medical device companies to enhance cybersecurity (U.S. Food and Drug Administration 2023).
While CVD has demonstrated effectiveness in traditional software security, its direct application to machine learning (ML) systems faces unique challenges. ML diverges from traditional software in two key aspects: 1) ML model issues must satisfy statistical validity thresholds, and 2) concerns related to trustworthiness and bias extend beyond the typical scope of security vulnerabilities (Office of Science and Technology Policy 2022). It is important to note that numerous critical bias reports discussed on social media and in the news may not align with the accepted definition of a vulnerability [^1] The Coordinated Vulnerabilities Enumeration program, managed by MITRE and central to the CVD ecosystem, defines a vulnerability as: A problem in software, firmware, hardware, or service component resulting from a weakness that can be exploited, causing a negative impact on the confidentiality, integrity, or availability of the affected component or components..
Upon recognizing this limitation in the conventionally accepted definition of a vulnerability, we propose to expand the terminology with a new proposed term - Flaw. We define a flaw as any unexpected model behavior that is outside of the defined intent and scope of the model design.
In response to the unique challenges posed by machine learning (ML) systems, we advocate for the adaptation of the Coordinated Vulnerability Disclosure (CVD) process into a dedicated ”Coordinated Flaw Disclosure” (CFD) framework. Tailored to ML’s distinctive properties, CFD aims to formalize the recognition of valid issues in ML models through an adjudication process. This process seeks to strike a balance between the interests of vendors and the broader community, ultimately enhancing public trust by curating disclosed flaws into actionable reports.
In the subsequent sections of the paper, we delve into specific aspects of this proposal. Section 2 provides essential background information, drawing a comparison between periodic audits and ad hoc harm reporting, while also exploring emerging participatory auditing practices. Section 3 examines evolving trends in disclosure practices within ML systems and identifies their shortcomings. Section 4 outlines the proposed changes within the CFD framework to improve disclosure rigor. Section 5 explores the practical implementation of CFD, proposing the utilization of model cards to define scope and intent. The paper concludes by emphasizing the potential of CFD to fortify AI accountability.
2 Background
In the following, we describe the background of our work. We will use the terminology introduced in the glossary in Table 1.
Concept Description
Vulnerability A problem in software, firmware, hardware, or service component resulting from a weakness that can be exploited, causing a negative impact on the confidentiality, integrity, or availability of the affected component or components.
Flaw Any unexpected model behavior that is outside of the defined intent and scope of the model design.
Coordinated Vulnerability Disclosure (CVD) Process in which vulnerabilities will only be disclosed to the public after reporting to the software vendor and giving them sufficient time to address the vulnerability.
Common Vulnerabilities and Exposures (CVE) A system managed by the MITRE Corporation under the National Cybersecurity FFRDC, offering a standard method for cataloging publicly known information security vulnerabilities and exposures.
Coordinated Flaws Disclosure (CFD) The framework proposed in this paper to address risks derived from flaws in ML models, analogous to CVD.
Common Flaws and Exposures (CFE) Analogous to CVE, the proposed system to report flaws of ML models.
Model Cards First proposed by Mitchell et al. (2019), model cards standardize the documentation of ML models, including the intended use cases of a reported model.
Table 1: Glossary.
2.1 Coordinated Disclosure Conventions in Cybersecurity
The cybersecurity community has cultivated a robust, 30-year history of well-established vulnerability disclosure practices, prominently featuring Coordinated Vulnerability Disclosure (CVD). This practice, acknowledged as a cornerstone of cybersecurity, facilitates collaboration between software vendors and vulnerability reporters to identify and address security flaws. Given the impossibility of eliminating all vulnerabilities during production, CVD establishes a crucial post-publication process for addressing these issues. Organizations, recognizing the limitations of in-house efforts, benefit from external parties and security researchers who disclose vulnerabilities through the CVD process. This collaboration allows affected organizations and the public to proactively address vulnerabilities, preempting exploitation by malicious actors. Moreover, the CVD process acts as a shield for those reporting vulnerabilities, demonstrating their good faith and avoiding legal conflicts (Cybersecurity Coalition 2019).
CVD’s significance extends to government policies, guidance, and regulations. United States civilian government agencies (Cybersecurity and Infrastructure Security Agency 2020) and numerous federal IT contractors (116th Congress 2020) are mandated to adopt vulnerability disclosure policies. The Food and Drug Administration incorporates CVD into its pre- and post-market guidance for medical device security (U.S. Food and Drug Administration 2023). Key standards and guidelines from the National Institute of Standards and Technology (NIST), such as the NIST Cybersecurity Framework (National Institute of Standards and Technology 2018), endorse CVD as a core practice. The EU’s NIS 2 Directive (European Parliament and Council 2022) requires critical infrastructure entities to establish CVD processes, and the recently passed EU Cyber Resilience Act (EUC 2023) mandates software manufacturers to adopt CVD processes. The European Union Agency for Cybersecurity further promotes national CVD legislation for its member states (European Union Agency for Cybersecurity (2022) ENISA). Additionally, CVD features prominently in international standards and industry best practices on security.
CVD operates within a broader framework of programs that identify, index, and communicate security vulnerabilities, including the National Vulnerability Database, CVE system, and Common Vulnerability Scoring System (CVSS). These programs support diverse security operations such as vulnerability management, risk assessments, penetration testing, patch management, and threat intelligence (Cybersecurity Coalition 2019). CVE Numbering Authorities (CNA) are organizations external to MITRE that are authorized to issue CVEs within their scope, but MITRE acts as a ”root” that handles disputes. For example vulnerabilities in open source projects managed by the Apache Software Foundation will be sent to them first and then, if needed, appealed to MITRE (Apache 2024). In the EU, the European Cybersecurity Scheme on Common Criteria (EUCC) is currently developed to cover information and communications technology (ICT) based on the international Common Criteria scheme (European Union Agency for Cybersecurity (2024) ENISA).
In the realm of machine learning systems, MITRE recognizes two ML CVEs: CVE-2019-20634 (CVE 2020) and CVE-2023-29374 (CVE 2023b). While CVE-2023-29374 involves a remote code execution with an LLM exploit, the vulnerability detailed by CVE-2019-20634 in Proofpoint’s system is less straightforward from a security perspective, since creating an exploitable copycat ML model requires a ton of positive and negative examples and their classification scores. Skylight Cyber’s discovery of a vulnerability in Cylance’s malware detection system, although awarded a Vulnerability Note by CERT, did not receive a CVE (VU4 2019). Evaluating the impact of these ML vulnerabilities is complex, and the existing coordinated disclosure infrastructure lacks precedence for such cases. Moreover, socioeconomic flaws in ML models (Solaiman et al. 2023), not categorized as vulnerabilities (see Table 1), also demand reporting.
Recognizing the importance of CVD to security, there is a temptation to extend these structures to encompass the disclosure of ML issues. However, existing CVD processes, without substantial modifications, prove insufficient to address the diverse range of ML issues. The disparities between ML models and traditional software necessitate a tailored approach. Nonetheless, the fundamental value of CVD in cybersecurity underscores the urgency of establishing a CVD-like process for ML models.
2.2 Differences between Traditional Software and ML in Disclosure
Traditional software vulnerabilities, such as buffer overflows, SQL injections, cross-site scripting, etc., typically arise from errors or weaknesses in the code, allowing attackers to compromise key properties like confidentiality, integrity, and availability.
In contrast to traditional software vulnerabilities with clear remediation steps, addressing ML model issues may involve retraining the model, modifying its architecture, or redefining its purpose. Furthermore, due to the probabilistic and complex nature of ML models, judgments are often difficult as the models lack distinct exploited and non-exploited states (Barbierato and Gatti 2024).
Most consumer ML models are deployed as closed-source APIs (Allison 2023), posing challenges in tracking and indexing vulnerabilities, especially in Software as a Service (SaaS) products. Versioning is not apparent to end users, and after remediation, the problem is not reproducible. Although these issues are recognized (Nichols 2021), CVEs for SaaS deployments are rare but have been issued (CVE 2023a). Continuous retraining and development add complexity to many AI systems.
These fundamental differences render existing vulnerability disclosure frameworks less applicable to ML model issues. Checking if an issue violates model integrity or intended functionality is less straightforward than identifying and replicating a software vulnerability due to ML models’ lack of causal explanations about achieved predictions (Barbierato and Gatti 2024). In the following subsections, we delve deeper into the unique challenges posed by ML systems that existing vulnerability management frameworks cannot fully address.
ML Models are Expected to Make Mistakes
At its essence, the distinction in reporting known issues between traditional and ML-based systems lies in the expectation that ML systems will make mistakes. While these isolated incidents may be inconvenient for users, they are anticipated, and both the vendor and users should have processes in place to address this inherent limitation of machine learning. It is crucial to consider mistakes collectively, incorporating them fairly into an unbiased dataset from a statistical standpoint before deeming them as issues requiring attention.
Users anticipate that the spam model safeguarding their email will make occasional errors. A report stating that it failed to accurately classify an email provides limited utility to users and vendors, often necessitating only a minor update to the training dataset. Even when scaled up with thousands of missed emails, it may not significantly impact processes or code. The report becomes valuable when it outlines an easily executable process to circumvent the spam filter. Detailed evasion process reports may introduce operational requirements that extend beyond dataset updates and scheduled retraining, impacting the model’s functionality.
However, there are instances that are entirely unacceptable to the model developer, elevating them instantly from incidents to issues. This spectrum ranges from rare, not-even-once incidents that qualify as issues to billions of data points collectively pushing the needle into issue territory. Identifying where on this spectrum an incident falls necessitates a case-by-case evaluation and is subject to various potential failures.
Intent of ML Model is Hard to Quantify and Test
Navigating ethical questions in disclosure adds complexity. Consider a model claiming demographic unbiasedness, verified by an independent auditor. If later discovered to be biased against an unexamined group, disclosure becomes crucial. Members of the affected group and users must adapt their use with this knowledge. The vendor deserves an opportunity to address and rectify the issue. Simultaneously, the independent auditor should enhance coverage for future models.
Yet, addressing ethical challenges, like bias, poses a drawback. Reports on such matters are challenging to assess. A report for a purportedly unbiased model may itself be biased, stemming from model mistakes or other biases. Distinguishing a biased report from legitimate ones becomes a challenge for external observers. While sampling biases can be refuted with a comprehensive model performance view, determining legitimacy may be impossible. Users need not alter behavior if the report is illegitimate, provided they already account for known error rates. The vendor and auditor need minimal behavioral changes, primarily preparing statements if the submitter opts to publish the report.
Ethical and Safety Questions are Hard to Judge
A pertinent example illustrating the mismatch with traditional vulnerability disclosure involves Facebook’s site integrity models (Noorshams, Verma, and Hofleitner 2020). These models address issues like hate speech, spam, and harassment, impacting the system. The intricate site integrity rulebook governs the trust and safety system, encompassing machine learning models and human reviewers. The complex objective function covers socioeconomic factors for billions across diverse social groups.
Evaluating if a model’s behavior violates this intricate code, distinguishing it from an understandable mistake in the moderation system, presents challenges. Submitters may create reports that, while seeming damning, fail to reveal a flaw in the moderation system. Vendors might reject valid reports on unverifiable statistical grounds, asserting the system operates within parameters. Both scenarios hinder the public – one erodes trust in an effective system, the other avoids fixing real issues.
2.3 Absence of Unified Product or Weakness Enumeration
Since the inception of the Common Vulnerabilities and Exposures (CVE) framework, an entire ecosystem has developed around it. In the National Vulnerability Database, a comprehensive CVE entry is now expected to encompass three key components:
  1. 1.
    Common Platform Enumeration (CPE) – Serving as a standardized naming scheme, CPE facilitates the identification of products that may contain vulnerabilities (Radack 2010).
  2. 2.
    Common Weakness Enumeration (CWE) – A collaborative initiative that meticulously enumerates weaknesses found in various products (MITRE 2024a).
  3. 3.
    Common Vulnerability Scoring System (CVSS) – An objective scoring system that quantifies the severity of a CVE through a numerical score (FIRST 2024).
These components, much like the CVE itself, significantly streamline the process of communicating and managing vulnerabilities within the cybersecurity landscape. The CPE allows organizations to quickly determine if their systems are susceptible to a particular CVE. Meanwhile, the CWE acts as a universal language for understanding the root causes of CVEs, thereby enhancing remediation efforts. Formally established in the mid-2000s, years after the CVE’s introduction in 1999, these components have become integral to the overall CVE ecosystem.
In the realm of machine learning ethics, several independent efforts have emerged to develop a classification system for potential harms caused by ML models (Weidinger et al. 2022; Shelby et al. 2023; Newman 2023; Graziani et al. 2023). Concurrently, within the domain of ML security, initiatives like the OWASP Top 10 for Large Language Models (LLMs) (OWASP 2021a) and ATLAS (MITRE 2024b) have gained prominence. Despite these strides, the ML community has not yet reached a consensus on a universally accepted structured harm reporting taxonomy. This challenge, however, reflects a familiar trajectory seen in the evolution of the CWE project. Drawing inspiration from contributions such as the OWASP Top 10 project (OWASP 2021b) and notable endeavors like PLOVER (Christey 2005), the CWE project successfully integrated diverse initiatives within focused, multi-stakeholder communities. This historical precedent suggests that the establishment of a standardized taxonomy in the ML community is indeed achievable.
3 Disclosure Practices in Machine Learning Systems: Evolving Trends
In this section, we examine the evolving trends of disclosure practices for machine learning systems and work that remains to be done for harmonization of these conventions. We break our analysis down into two axes – objective security and privacy flaw disclosures, and subjective ethical flaw disclosures, such as societal bias.
3.1 Security Disclosures for ML Systems
ML-based systems can be subject to three main types of disclosures from users and researchers: security vulnerability reports, model design failures, and error reports. In the following discussion, we examine examples under each category. Among these reports, some highlight actionable issues warranting recognition by external authorities, while others delve into details that may not demand significant attention.
Infrastructure Reports
These reports highlight security vulnerabilities in ML models that are independent of the model’s performance. These are not usually model vulnerability reports, but are in components around the model. They often stem from traditional vulnerabilities in the components, in which case they should be referred to the Common Vulnerabilities and Exposures Numbering Authority (CNA) responsible for that specific component.
CVE reports on Numpy (CVE 2019a), Tensorflow (CVE 2019b), and Pytorch (CVE 2023c) fall into this category. The OpenAI bug bounty on Bugcrowd also addresses infrastructure flaws or vulnerabilities, as issues in the AI itself are considered out of scope (OpenAI 2023). Numerous examples exist of vulnerabilities that model inference pipelines might face, as the machine learning community often prioritizes reproducibility of experiments over code security. Efforts to address this concern are emerging, such as SafeTensors by Hugging Face (Hugging Face 2022).
An instance of a traditional security vulnerability caused by an ML model is DeepSloth (Hong et al. 2020), presenting a Denial Of Service attack similar to SlowLoris (Cloudflare 2024). This type of attack could be executed by an external entity on a Large Language Model (LLM) using early exit or multi-exit strategies to minimize costs. Consequently, this may lead to overprovisioning, reaching capacity, or system failure. Several model optimization techniques used by LLMs in production could be potential targets for such attacks.
A novel vulnerability is ProofPudding (MITRE 2019), the first CVE issued for an issue caused by a machine learning model. Model stealing, though concerning, demands numerous queries, especially for large models with complex behaviors. The risk is significantly reduced when the API provides excessive information. Proofpoint returned the exact output to attackers, minimizing the number of queries needed. However, they failed to implement rate limits on the information. The mitigation applied by ProofPoint was to limit the information their API provides, a practice theoretically possible with explainable AI techniques like Shapley values (Oksuz, Halimi, and Ayday 2023). The implementation of rate limits and limited information in responses represents best practices in ML security, and we anticipate more reports of this nature in the future.
Model Design Failure Reports
Model design failures are typically identified by manipulating samples to induce changes in the output. While they may resemble security vulnerabilities, unlike infrastructure reports, these vulnerabilities rely on modifying the output of a deployed machine learning model. They are not addressable simply improving the underlying dataset (Hooker 2021) and are shared among all models of a similar type. The complexity of these issues may render them unsolvable with current technology, leading to reports of this kind being considered out of scope or low priority.
The first documented vulnerability of this type in a production system was a “Bayesian Poisoning Attack” reported in 2003 on a bulletin board about PoPFile (The MIT 2004 Spam Conference ), and later on Paul Graham’s Blog (Graham 2003). This easily reproducible bypass did not affect the content of the spam but was valuable due to its focus on the method. Another model exhibiting similar behavior was the Cylance malware detector (Cyber 2019). Additional examples relevant to security practitioners come from the Malware Detection Evasion Competition, designed to elicit reports of these types (Anderson 2019). The competition rules included banning known bypasses like packing and awarded prizes for the most novel and interesting techniques (Anderson and Balazs 2020).
One frequently discussed model design failure is Adversarial Examples, which many machine learning security researchers consider low priority at best (Szegedy et al. 2014). These examples are often delicate, requiring a high level of technical expertise and control over the underlying system. Adversarial examples frequently do not survive jpeg compression (Aydemir, Temizel, and Temizel 2018) or a simple blur & repair (Carlini et al. 2022). Although one cited use for this attack is in self-driving cars, such scenarios usually demand extensive compromise of the car’s vision system. Adversarial patches represent a potential attack path (Brown et al. 2018) but are challenging to execute (Athalye et al. 2018). Marines tasked with bypassing an ML model monitoring CCTV footage succeeded with a low-complexity attack by getting into a large box (Scharre 2023b). In most scenarios, lower-hanging fruit for attackers is more appealing (Apruzzese et al. 2022), often resulting in these issues being considered out of scope or low priority for most ML systems, as vendors prioritize other pressing concerns.
Error Reports
These cases involve issues that could stem from the expected statistical error rate demonstrated by the model during testing or systematic errors that are rectifiable through changes to the training data or modifications in how the model is trained. Such reports are ubiquitous, and vendors often incorporate telemetry and remediation tools into their systems for data gathering.
  • One-off Errors: These errors are an anticipated aspect of the model’s functioning. ML models seldom achieve 100% accuracy, and these errors are within the normal error rate and comply with the Service Level Agreement (SLA) when considered in conjunction with the rest of the query data.
  • Systematic Errors: These errors arise from unintended statistical biases in the model, leading to failure in producing output aligned with the intent of the model in one or more categories.
One-off errors are commonplace, such as spam ending up in the main inbox or legitimate emails being marked as spam. Actions like moving a legitimate email out of spam or reporting spam contribute to telemetry, enhancing subsequent model datasets. While these errors are typically perceived as expected inconveniences in using machine learning systems, they become reportable if they are egregious and cause harm. An extreme case is Google’s image classification system mislabeling a black person as a gorilla, which should be publicly reported using a coordinated disclosure process even if it occurs once (Scharre 2023a).
Systematic errors pose greater challenges in diagnosis, often originating from statistical biases in the dataset. These biases may result from socioeconomic forces, poor collection techniques, or unforeseen reasons during the model production, leading to harm. A classic example often shared with new machine learning students involves a tank classifier learning to recognize tanks based on the sky. British tanks were photographed with gray skies, and Russian tanks with blue skies, causing the classifier to learn the easier task of sky classification. While this specific story may be an urban legend, it serves as a cautionary tale for students (Branwen 2023). Another recent example of a systematic error is the “divergence attack” (Nasr et al. 2023) demonstrated by Google researchers on ChatGPT – when ChatGPT was asked to repeat a random word like “poem” infinitely, it started leaking private training data.
3.2 Disclosure of Ethical Issues in ML Systems
Despite calls for audits from academics, lawmakers and activists alike (Raji et al. 2022; Fjeld et al. 2020; Wilkinson et al. 2023), the identification and disclosure of ethical issues in machine learning (ML) systems have, for the past decade, progressed in a somewhat ad hoc and reactive manner. Unlike traditional software systems, which follow established practices for coordinated vulnerability disclosure, the process of surfacing problems with ML models has been inconsistent at best. Due to the lack of a mandated reporting mechanism like the one we outline in this paper, harms caused by flawed ML systems are often only discovered after the damage is done (Scharre 2023a). Few proactive, systematic mechanisms exist for auditing these complex sociotechnical systems, identifying risks pre-deployment, and enabling responsible disclosure in a manner that balances organizational and community interests. However, promising approaches, such as cooperative audits (Wilson et al. 2021), participatory audits (Dennler et al. 2023) and AI red teaming (OpenAI 2023; Ganguli et al. 2022), are beginning to take shape as the stakes surrounding ML systems continue to grow. This section traces the disclosure of ethical issues in machine learning, covering its history and trends.
Reporting has primarily relied on affected communities facing issues and then raising complaints through social media campaigns, media coverage, lawsuits, and other public pressure tactics after the fact. There have been few systematic mechanisms for auditing and identifying model issues proactively during development phases or continuously monitoring systems post-deployment. Prominent examples include recidivism prediction algorithms like COMPAS exhibiting significant racial bias against Black defendants, only uncovered through investigative journalism efforts (ProPublica 2016). Facial analysis systems misclassifying people of color often go unnoticed until public backlash erupts on social platforms (BBC News 2015). Harmful biases in language models generating toxic outputs against marginalized groups are frequently observed only due to vigilance within directly impacted communities (Cosmos Magazine 2023; Abid, Farooqi, and Zou 2021).
The lack of established processes enabling coordinated discovery and disclosure of ethical issues in machine learning and the closed-source nature of many AI systems means users are unaware how long problems fester before coming to light. The reliance on reactive reporting by harmed parties has made accountability an exception rather than the norm for ML systems. Researchers and practitioners alike have highlighted a pressing need to complement one-off investigations with proactive, community-centered mechanisms for responsible disclosures. A key way to enable decentralized discovery and disclosure of ML ethical flaws include providing legal protections for independent ML issue researchers, in a similar fashion to existing legal protections for good faith security research (Hacking Policy Council 2023).
3.3 Participatory Approaches for Issue Discovery
In response to growing public scrutiny around AI ethics and recognition of the severe limitations of ad hoc reporting for machine learning systems, some organizations have recently started experimenting with participatory approaches to systematically discover model issues.
Bias Bounty Programs
Bias bounty programs have emerged as one mechanism enabling security researchers and external auditors to uncover harmful biases and other flaws in ML models. These programs incentivize crowd-sourced identification of algorithmic problems by providing monetary rewards and public recognition for impactful findings. Some prominent early efforts in this space include Twitter’s algorithmic cropping bounty after concerns over racial bias (Yee, Tantipongpipat, and Mishra 2021) and QueerInAI’s queer bias bounty (Dennler et al. 2023) targeting intersectional issues, but AI bug bounties are slowly becoming commonplace in big companies as well (Google 2023; Microsoft 2023).
Community Driven Audits
Additionally, community-driven audits (Cattell, Chowdhury, and Carson 2023; The New York Times 2023), petitions (CAIR 2023), and pressure campaigns (Bloomberg 2023) have proven effective for advocacy groups to highlight urgent ML issues and hold organizations accountable. Truly participatory approaches, however, require extensive investments towards inclusion to lower barriers for participation by directly affected groups (Dennler et al. 2023). They also necessitate extensive transparency into model development and deployment choices to foster informed, constructive feedback. In addition, these approaches presently tend to have the authorization of the ML model owners, which the ML model owners could withhold, rescind, or condition to silence potentially damaging disclosures under the threat of a lawsuit (Friedler et al. 2023). Depending on ML model owners for authorization to research the ML models can ultimately undermine researchers’ independence.
Crowdsourced AI Harm Databases
There are some nascent efforts towards standardized documentation of AI harms. Taxonomies such as MITRE’s ATLAS [^2] https://atlas.mitre.org/ and databases such as the AI Vulnerability Database (AVID) [^3] https://avidml.org/ managed by AI Risk and Vulnerability Alliance (ARVA), and the AI Incident Database (AIID) [^4] https://incidentdatabase.ai/ managed by the Responsible AI Collaborative are the most well known efforts in the field. AIID, well-established, collects news stories about AI incidents causing harm, while AVID assesses several open-source models and posts their findings on their database. Although useful, these initiatives lack coordinated disclosure features and mechanisms for vendor rebuttal. There are bounty programs like Google’s and Microsoft’s that demonstrate an embrace of disclosure at large model vendors (Google 2023; Microsoft 2023).
AI Red Teaming
AI red teaming has emerged as one of the most promising recent developments in surfacing ethical issues in machine learning before model launches (OpenAI 2023; Ganguli et al. 2022). It refers to an organization commissioning a diverse group of outside domain experts to rigorously test and attempt to attack their ML systems, simulating the behaviors of real-world malicious actors (Carey and Jin 2019; The White House 2023b).
The goals of AI red teaming exercises are to pressure test models, infrastructures, and deployment environments in adversarial ways, surface unknown issues or vulnerabilities that may have been overlooked during in-house testing, provide external validation of internal safety claims, and gain new perspectives for strengthening systems before public release (Selbst 2021; Moss et al. 2021; Carlini et al. 2023; Zou et al. 2023). Effective red teams require careful screening for a diverse mix of professional backgrounds, ideologies, regional representations, and skill sets.
While promising, these community-centered mechanisms remain the exception rather than the norm. Much work lies ahead in formalizing scalable, sustainable, and equitable participatory practices for ML issue discovery and disclosures.
4 Proposed Changes to Recognition of a Disclosure
In traditional coordinated disclosure for software vulnerabilities, if a vendor rejects acknowledging a reported issue, the only recourse for the submitter is to publicly disclose the problem and advocate for recognition. While this can be effective in drawing attention to and resolving vulnerabilities, the dynamics change when dealing with ML-based systems.
ML disclosures often involve ethical considerations, particularly regarding political speech. The challenge lies in establishing a process that fairly evaluates submissions, distinguishing substantiated claims from unsupported ones aiming to serve the submitter’s agenda. Accepting ethical issue reports based on sampling bias can erode public trust in the vendor, while rejecting valid issues supported by statistically sound data damages both public trust in the vendor and the entity responsible for the flawed rejection.
To address these challenges, we propose a Coordinated Flaw Disclosure (CFD) process that incorporates:
  1. 1.
    Extension of Model Cards: Model Cards, traditionally used for complete systems, will be extended to provide detailed documentation of intent and scope.
  2. 2.
    Trusted Independent Adjudication Panel: A panel of independent adjudicators will be established to mediate disputes between the submitter and vendor when seeking official recognition.
  3. 3.
    Automated Verification through Submission Extension: An extension of the submission concept will enable automatic verification of the potential issue’s existence, enhancing the efficiency of the recognition process.
The main difference between this and the current CVE process is that it is based on the specific model or system card, not a universal definition of vulnerability. Bug bounty programs usually have a “scope” document associated with each vendor that dictates what is and is not reportable. The CVE process does not have this pre-defined scopes for a system; the definition of the vulnerability and community norms manage it. Vendors may argue for a lower CVSS score from MITRE if they believe the vulnerability is not relevant. However, there are flaws that are reputable in one model but not in another due to the creator’s intent. Llama2 has an extensive model card (Meta 2023), but Mistral is meant to be a base model that is fine-tuned and embedded in another system, and the model card reflects that (AI 2023). Mistral’s card is quite sparse and incomplete for a process such as this.
The main component is the trusted independent adjudication panel to govern the disputes that will arise between the submitter and vendor. This is a key part of the official CVD processes, as demonstrated by the “root” in the CVE Numbering Authority (CNA) rules for 3rd party CVE enumeration (Board 2020), the Bugcrowd appeals process (BugCrowd 2024), and the HackerOne Mediation process (Hackerone 2024). This will provide an outlet that fairly hears the submitter but also the vendor. If there’s sensitive proprietary information from the vendor that refutes the issue, the adjudicator can review it and resolve the dispute without a public fight. However, if this data does not truly dispute the submitter, it can award the issue and thus start the mitigation process at the vendor.
We hope that this process will enable vendors to create safer and more secure models, and the public to understand them better. We expect the discovered issues that arise in the course of a model system’s development to impact its Model Card to make it more precise, thus enabling everyone to better understand the capabilities of the system.
4.1 Complete Model/System Card
The vendor needs to supply a model card for the complete system that includes intent and scope. These are related; intent covers what the system is meant to do, and scope limits the issues people can look for. This is already called for in the original paper (Mitchell et al. 2019) and implemented in many cards, including the Llama2 card (Meta 2023). Intent includes a mission statement for the system and the efficacy measures the model is tied to. These will evolve over time as the model is better understood in its environment. Scope is focused on known issues that the model has that are either in the process of being addressed or cannot be resolved in a reasonable timeframe. These irresolvable flaws come in two flavors: fundamental flaws in the system that can never be resolved and known flaws that are being worked on but might take a few years.
Essentially, this breaks up the idea of “scope” from the security community into two terms. This is needed as the idea of fallibility is built into our use of the term “intent,” but “in-scope” for security professionals means any vulnerabilities found in that scope. In security, the intent is often understood by the product description and community understanding. SQL injections are not a flaw in databases unless that database has a permissions scheme that the query was able to violate. As identical models, in terms of the input and output formats, can have wildly different intents, we believe that an intent statement will clear the air and reduce the amount of scope details needed. The Microsoft bounty program makes this distinction clear for LLMs, excluding prompt injections unless it can get privileged information from other users.
This model card is a complete system card. This system can contain known vulnerable models and additional models that are meant to mitigate those vulnerabilities. For example, using a diffusion model with an image net classifier might be secure against adversarial examples, and the model intent and scope can clearly state this as a system property where it is not a property of the individual model. This would make adversarial examples “in-scope” and reportable. Another example, for an LLM, you can deploy a hate speech model with the LLM that blocks hate speech like with NVIDIA Guardrails (Rebedea et al. 2023). The LLM model card should include “this potentially produces hate speech,” but the system can say “this does not produce hate speech.” Another example is Dalle2 inserting demographic descriptors into prompts to increase the diversity of the produced output (OpenAI 2022).
We now further elaborate on what “Intent” and “Scope” mean in the context of our proposed framework.
Intent
To make “intent” actionable, the system card should include a mission statement for the model system, and as many precise efficacy measures as possible.
Examples of Intent Mission statements:
  1. 1.
    From a text prompt, we produce images that are safe for work, safe for children, and free of demographic bias.
  2. 2.
    This malware model classifies Portable Executable (PE) binaries into malicious or benign. We consider Potentially Unwanted Applications, such as adware, to be malicious.
Examples of Intent measures:
  1. 1.
    We reproduce images of professionals in particular careers at the demographic rates of the US in this Text to Image model.
  2. 2.
    We have a 0.001 False Positive Rate and a 0.1 False Negative Rate on each customer’s cumulative data for this malware model, if that customer has over 100 endpoints.
Scope
The main purpose of Scope is to exclude known issues that the vendor has no intent or ability to resolve in the next year. There are intents, like Mistral’s release of a base model for use in a larger system, that essentially exclude many reportable flaws. The two scope documents from the Google and Microsoft AI Bounty programs provide good examples of in and out-of-scope statements.
  1. 1.
    L2 norm adversarial examples are a known vulnerability in this image model; do not use for sensitive applications.
  2. 2.
    This is an LLM with no protections; prompt injections are out of scope.
  3. 3.
    Binary packing is a known vulnerability of this malware detection model, and we have no plans to build mitigations.
4.2 Adjudication Process
The vendor should not have sole authority over the recognition of an issue, especially when they are the only entity that can verify the issue’s existence. Additionally, reports should be carefully vetted to ensure that the issue truly exists, is a valid violation of the model card, and is not a result of sampling bias in the provided report samples. We acknowledge that this is a significant deviation from the CNA process used to grant vendor CVEs and anticipate that it will be a topic of discussion.
The primary function of this committee is to distinguish between issues that involve valid violations of the model’s intent and scope and those that do not. One challenge arises when an unbiased model used in hiring or image generation may not precisely align with a demographic split. If the model card does not allow for an error in this measurement, it would be considered a violation in the strictest sense. However, if the error is within industry standards, a report detailing this may not be officially recognized.
There are instances where the decisions made by the model may violate one of the stated intents of the model but align with another. Achieving complete political unbiasedness may be challenging, especially if one political group violates another policy far more frequently than another group.
4.3 Automatic Verification
Errors in the model are going to be a large source of reports to any database of model issues. Reporting these to the vendor through a defined reporting system, a telemetry system for example, is preferable for the vendor for several reasons, but most importantly it enables easy integration into the model’s training and testing dataset. It also verifies that the model actually has the behavior and it’s reproducible, which may change day-to-day with model updates and continuous training. An outside report requires labor from the vendor to verify the behavior and add the input/output pair into the dataset.
Almost all useful reports listed in the background section include samples that demonstrate the behavior. Any system that aids in Coordinated Flaw Disclosure with the vendor should include support for automatic external verification and automatic integration in the vendors internal testing dataset. This significantly lowers the costs to the vendor to potentially remediate the issue, and provides a means of checking that the issue stays closed. If the issue is accepted the system should retain the data used to support the report and periodically re-verify the issue. If it is resolved by the vendor in one model, a subsequent model may regress and re-open the issue.
Figure 1: A potential template for a CFE report by a user.
5 Issue Reporting Governance
In the following, we describe the process of submitting a CFD. The implementation of the process depends on the vendors, however, their clear advantage of formalizing the process is similar to the adoption of the previous CVD process; vulnerabilities are explored before they can be exploited. Where CFD-related issues are identified, we can assume that it is in the best interests of the vendors to be notified about those to be able to resolve them in a timely manner. Our proposed CFE issuance process (Section 5.2) takes this into account, and proposes a process to resolve potential issues between reporters and vendors.
5.1 Submitter Interaction
Users identify potential issues believed to violate the scope and create a submission with:
  1. 1.
    The sample set (even a single sample is acceptable)
  2. 2.
    The output(s) the model should not generate.
The report argues that the sample output pairs violate the scope. After submission, the report is sent to the vendor for acceptance or rejection. If rejected, the user can appeal to the Adjudicators.
We propose a reporting template in Figure 1. The submitter should provide information about the vendor, system, and the person reporting the issue. The description of the uncovered flaw should include whether this is a infrastructure or model design failure report, and whether it is a one-off error, or a systematic error. For more details on the error types, see Section 3.1. To give the vendor context about how to address the issue and a possibility to prioritize the report, the report should refer to the model card of the model. Finally, the template should have the possibility to submit sample set and according output, depending on the size either as a table or as structured in a common data format.
5.2 CFE Issuance Process
An overview of the CFE issuance process can be seen in Figure 2. Upon submission of a flaw, the following steps are followed:
  1. 1.
    Triage: The vendor or an authorized person reviews the report to ensure the sample and output pairs satisfy the report.
    • Note: If not, the submission is rejected on technical grounds.
  2. 2.
    Review: If the report and sample/output pairs align, the vendor determines if the report is within Scope and violates the model Intent.
    • Note: If not, the submission is rejected as either out-of-scope or in-intent.
    • Caveat: If the data supports the report but is incomplete or flawed, the vendor can reject on statistical grounds or request additional data.
  3. 3.
    Adjudication: If rejected for any reason, the Submitter can appeal to the adjudicator panel.
The vendor may reject the submission on statistical grounds to the adjudicator panel. If the submitter-supplied sample/output pairs are statistically biased, and an unbiased sample set would not show a violation of the model card, the adjudicator panel may request supporting data from the vendor to validate the rejection. The adjudicator and vendor can jointly assess whether the vendor-supplied data is too sensitive for the submitter and decide on further steps. An example of this cooperation is once again the “divergence attack” paper by Google (Nasr et al. 2023). The authors followed a standard 90 day responsible disclosure practice [^5] https://not-just-memorization.github.io/extracting-training-data-from-chatgpt.html#responsible-disclosure in their private disclosure of the bug to OpenAI and publicly reported the bug only after the 90 days were up.
5.3 Adjudicator Panel
Current policy proposals focus on providing audits by external experts (Hartmann et al. 2024). These audits have a very wide range of tasks and possible restrictions to apply. In the CFD process, these already planned panels could also act as adjudicator panels in the CFD processes. It would give the audits a starting point about contested issues, and could be a starting point for the implementation of the CFD process. It is important to emphasise that we expect the vast majority of CFD to be resolved between the reporters and AI providers directly as it is the case with the CVE processes (Red Hat 2022; MITRE Community Board 2014), keeping the workload of the adjudicator panel light.
6 Conclusion
In this paper, we have proposed a Coordinated Flaw Disclosure (CFD) framework tailored specifically for machine learning and artificial intelligence systems. Our framework addresses the nuanced challenges posed by algorithmic flaws, emphasizing the importance of transparent intent and scope definitions through an extension of Model Cards for complete systems. By incorporating a trusted independent adjudication panel, we aim to facilitate fair dispute resolution, effectively distinguishing valid violations from issues rooted in sampling biases. The introduction of an automatic verification mechanism streamlines reporting processes, enhancing efficiency and ensuring issue reproducibility. With a structured issue reporting governance model encompassing triage, review, and adjudication, the CFD framework seeks to strike a balance between submitter and vendor interests, fostering trust and transparency. We believe that widespread adoption of this framework can contribute to bolstering public trust in AI systems by promoting clear processes and collaborative engagement between organizations and the broader community. As the field of machine learning evolves, the CFD framework provides a structured and adaptable approach to navigate the ever-changing landscape of algorithmic flaws, benefiting both vendors and the public alike.
Figure 2: Flowchart of the CFE issuance processes described in Section 5.2.
6.1 Limitations
Model Cards need Standards
A notable limitation in implementing a CFD process is the lack of standardized model cards. Existing system card formats exhibit significant variation, and many models utilized by vendors lack public model cards. Addressing this discrepancy is crucial for widespread CFD adoption. We echo calls from the NIST AI Risk Management Framework for the establishment of standards for model cards (Singh 2021), providing a necessary foundation for developers, especially those with commercial ambitions.
Challenges of Dynamic ML Datasets
Another challenge stems from the complex nature of models built on diverse datasets, each potentially requiring a data card (Pushkarna, Zaldivar, and Kjartansson 2022). Drawing a parallel from the software domain, the concept of a Software Bill of Materials (SBOM) (CISA 2021) has been instrumental in regulating version control for software components (Corman 2022), essential for avoiding CVEs. However, the absence of equivalent tools or practices for tracking Dataset cards poses a significant hurdle. Unlike the well-established open-source tools for determining software dependencies, tracking the datasets used in model training, which evolve over time, presents a more intricate problem. Recent incidents, such as the retraction of the LAION-5B dataset (Schuhmann et al. 2022) due to the presence of Child Sexual Abuse Material (CSAM) (LAION.ai 2023; Thiel 2023), underscore the pressing need for robust mechanisms to manage dataset-related challenges in the machine learning landscape.
Common Weaknesses Enumeration: Contextual Challenges
Introducing a Common Weaknesses Enumeration (CWE) system for machine learning introduces novel challenges. Even for vulnerabilities, the weakness enumeration is highly context-specific. For example, a system might be theoretically vulnerable to the ATLAS (MITRE 2024b) weakness “Prompt Injection,” but due to the application’s design, this may not be a legitimate concern. In traditional software, a memory weakness like a buffer overflow in a C program is consistently a concern, often leading to vulnerabilities allowing remote code execution. While context dependency is not novel for AI systems, the introduction of identical ML models for very different purposes adds a layer of complexity. Unlike most traditional software contexts, the documentation around the model becomes crucial, introducing a dependency not typically encountered in conventional software scenarios.
7 Research Ethics and Social Impact
Our interdisciplinary team, consisting of AI ethicists and security researchers, navigates the intersection of AI ethics and cybersecurity, drawing inspiration from established practices like the Common Vulnerabilities and Exposures (CVE) process. We recognize the ethical imperative of mitigating harms associated with vulnerability reporting, aligning with the broader goal of enhancing model transparency and promoting the development of safer AI systems.
7.1 Ethical Considerations
In adherence to ethical research principles and community norms, we disclose potential challenges and ethical considerations associated with our work. No human subject experiments were conducted, and therefore an IRB was not required.
7.2 Author Positionality
As authors with a foundation in privacy and cybersecurity, we have witnessed the positive outcomes of the Coordinated Vulnerability Disclosure (CVD) process for software projects, both small and large. This proposal advocates for extending the CVD process to the field of machine learning research, aiming to bridge the gap and introduce machine learning researchers to the benefits of a familiar and effective process. Our positionality as proponents of CVD in this context stems from our understanding of its commercial advantages and positive impact on software security.
7.3 Adverse Impact
In the majority of cases, the CVD process operates smoothly without adverse consequences. However, we acknowledge instances where a vendor may reject an issue without engagement, addressing it quietly without public acknowledgment. While such occurrences are limited, they emphasize the need for a framework that encourages responsible disclosure and engagement, ensuring transparency and accountability in the machine learning community.